OpenStudy (anonymous):

Write the repeating decimal 0.2666666... As a fraction in simplest form. It's our really just like 4/15 because that seems too easy

hero (hero):

let x = 0.2666666... 10x = 2.666666... 10x - x = 2.666666... - 0.2666666 9x = 2.4 90x = 24 x = 24/90 x = 4/15

hero (hero):

@msarahrileyc

OpenStudy (anonymous):

Oh thanks